# Tunable lower critical fractal dimension for a non-equilibrium phase transition

### v1

@@ Abstract
Our results were obtained by solving the many-mode mean-field dynamics and analyzing the emergence of critical slowing down as the size of the driving patterns increases. An open and important question concerns the fate of the lower critical dimension in the presence of strong quantum boson-boson interactions, beyond the many-mode mean-field regime, particularly in the hard-core boson limit. Three scenarios could take place: (i) the lower critical dimension is the same in the strongly correlated regime; (ii) the lower critical dimension could evolve continuously with interaction strength, or (iii) a discontinuous shift could occur, signaling the emergence of a quantum phase transition with new phases in the strongly quantum correlated regime. Exploring this limit remains a key direction for future research. Another promising avenue is to investigate how fractional dimension and driven-dissipative conditions influence other classes of non-equilibrium phase transitions. This could yield deeper insight into the fundamental role of spatial geometry and dissipation in shaping critical phenomena in classical and quantum open systems.
@@ Acknowledgements.
“Tunable lower critical fractal dimension for a non-equilibrium phase transition” June 25, 2025
@@
I The effect of detuning for a Sierpinski carpet
@@
In the main text, we characterize critical behavior in terms of the minimal asymptotic decay rate. Here, we further examine how the asymptotic decay rate depends on the driving intensity for different system sizes. In Fig. S.1(a)–(b), these quantities are shown for the Sierpinski carpet (first row in Fig. 1) with $\Delta=\gamma$, a regime where no phase transition occurs. The minimal asymptotic decay rate exhibits a clear minimum around a critical driving intensity $I\approx 2.2\gamma^{2}$, yet this minimum saturates to a finite value. Correspondingly, the density increases smoothly across all system sizes, deviating from the homogeneous single-mode mean-field prediction [4] (dashed grey curve) obtained by retaining only the homogeneous $k=0$ mode.
@@
Figure S.1: Response of the system to a Sierpinski carpet drive with fractal dimension $d_{H}=1.84$. (a) Asymptotic decay rate as a function of input intensity for detuning $\Delta=\gamma$, showing no phase transition. (b) Corresponding bosonic density as a function of input intensity, with a smooth increase. (c) For a larger detuning $\Delta=2.0\gamma$, the minimal asymptotic decay rate $\lambda_{\min}$ vanishes at the critical point $I_{c}\approx 7.1\gamma^{2}$, indicating a first-order phase transition. (d) At the same critical point, the average bosonic density exhibits a discontinuous jump.
@@
In Fig. S.1(c)–(d), we plot the same quantities for a larger detuning, $\Delta=2\gamma$. In this case, at a higher driving intensity $I_{c}\approx 7.1\gamma^{2}$, the minimal asymptotic decay rate $\lambda_{\min}$ collapses to numerical precision, and the density exhibits an abrupt switch from the lower to the upper branch of the bistability. These features indicate the occurrence of a first-order phase transition.
II Box-counting dimension
@@
To quantify the fractal dimension of the considered patterns, we employed the box-counting method [31]. The shape of each pattern is first extracted from a grayscale image (see, for example, Fig. 1) by applying a threshold above which points are considered part of the pattern [34]. This thresholding procedure delineates the fractal structure within the image, effectively separating the foreground (pattern) from the background and yielding a binary representation. In this binary image, pixels above the threshold are marked as belonging to the pattern, while those below are discarded. This representation serves as the basis for further spatial analysis.
@@
The box-counting method then divides the binary image into a grid of $s^{2}$ square boxes, each with edge length $b=L/s$, where $L$ is the linear size of the system. We count the number of boxes that intersect the pattern, thereby estimating the area $A$ of the fractal (in units of $b^{2}$). In principle, this area scales as $A\propto s^{d_{H}}$, where $d_{H}$ is the Hausdorff (fractal) dimension. The box-counting dimension is thus given by the slope of $\log A=f(\log s)$ obtained over different scales $s$. By performing a linear fit to this curve, we extract an estimate of $d_{H}$, and the corresponding uncertainty is reported in the main text as the error bar on the box-counting dimension.
The patterns in Fig. 1 are annotated with their respective box-counting dimensions $d_{H}$, which differ slightly from the theoretical values $d^{th}_{H}$ for perfect, infinitely self-similar fractals. For example, for the Sierpinski carpet (second column), $d^{th}_{H}=1.89$, and for the Sierpinski cross (third column), $d^{th}_{H}=1.46$. Nevertheless, even for these distorted and finite structures, we can determine a fractal dimension lying between 1 and 2.
@@
For patterns with fractal dimension equal to or lower than 1 (Fig. 1, last column), we used a slightly modified version of the box-counting method. We restricted the analysis to the horizontal dimension by fixing the vertical box size to the system size $L$, while the horizontal box length still scales as $b=L/s$. Instead of square boxes of size $b^{2}$, we thus used rectangular boxes of size $bL$. This allows for a consistent evaluation of fractal dimensions ranging from 0 to 2.
For integer-dimensional patterns (1D and 2D), we did not apply the box-counting method, as their Hausdorff dimensions are trivially $d_{H}=1$ and $d_{H}=2$, respectively. Nonetheless, one could apply box-counting to the patterns shown in the first and fourth columns of Fig. 1; such an analysis would yield values close to 1 or 2, but never exactly. Indeed, precisely capturing integer dimensions via box-counting is challenging due to inherent estimation errors.
III The spatial profiles of the studied fractals
@@
Figure S.2: Left panels: 2D square geometry ($d_{H}=2$) and several variations of the Sierpinski carpet with fractal dimensions in the range $21<d_{H}<$ 2. Right panels: Variations of the Sierpinski cross ($21<d_{H}<$ 2) in the first two rows, followed by the 1D strip ($d_{H}=1$) in the third row. The final row shows variations of the Cantor set with $d_{H}<1$. All patterns are computed with $\Delta=\gamma$, and the driving amplitudes are tuned to their respective critical points. The geometries and system sizes correspond to those used in Fig. 3, except for the last geometry on the left, which is excluded from the main text due to lack of convergence in the asymptotic decay rate for the considered spatial sizes.
@@
Figure S.2 shows the spatial profiles of the patterns analyzed in the main text (Fig. 3). In addition to the standard 2D and 1D reference patterns, we include several variations of the Sierpinski cross and carpet, as well as modified versions of the Cantor set with fractal dimension lower than one. One particular geometry (the last one on the left in Fig. S.2) is not shown in the main text, as the asymptotic decay rate did not converge in the thermodynamic limit within the range of system sizes we were able to simulate.
IV Dependence of the critical driving on the detuning
@@
In Fig. S.3, we extract the value of the driving intensity $I_{\min}$ such that $\lambda(I_{\min})=\lambda_{\min}$, for a range of detuning values. As a reminder, $\lambda(I)$ denotes the asymptotic decay rate of the observed pattern as a function of the driving intensity $I$, and $\lambda_{\min}$ is its minimum value. As discussed in the main text, this minimum signals the presence of critical behavior in the system. Plotting $I_{\min}$ as a function of detuning thus allows us to trace how the onset of criticality evolves under varying external conditions.
@@
Interestingly, the value of $I_{\min}$ closely matches the inflection point of the bistability curve derived from single-mode ($k=0$) mean-field solution for the homogeneous driving case. The inflection point corresponds to the value of the driving intensity at which the curvature of the response function changes sign—specifically, where the second derivative of the driving intensity with respect to the density $n=|\psi|^{2}$ vanishes:
$\frac{d^{2}I}{dn^{2}}=0.$ (S.1)
@@
$I_{\min}=\frac{2\Delta}{3g}\left(\frac{\Delta^{2}}{9}+\frac{\gamma^{2}}{4}\right),$
which is shown as the grey dashed line in Fig. S.3. The close agreement between the numerical values of $I_{\min}$ and the analytical inflection point provides a consistency check and further supports the interpretation of $\lambda_{\min}$ as a reliable marker of criticality.
Figure S.3: Evolution of the critical point $I_{\min}$ as a function of $\Delta$. A quadratic dependence is observed, consistent with estimates based on the single-mode mean-field description of the bistability curve. The inflection point of the mean-field curve (grey line) closely approximates the critical point across all geometries considered.
V Asymptotic decay rate and Bogoliubov instabilities
For a homogeneous driving, we can diagonalize the problem for fluctuations on top of the steady state and obtain a spectrum of the so-called dissipative Bogoliubov modes [4]. By analyzing the dispersion relation given in equation (S.2), we can interpret the phase transition as the emergence of an unstable mode.
$\omega_{\mathrm{Bog}}(\mathbf{k})=\pm\left[\left(-\Delta+\frac{\hbar\mathbf{k}^{2}}{2m}+2gn\right)^{2}-\left(gn\right)^{2}\right]^{1/2}-i\frac{\gamma}{2}$ (S.2)
@@ V Asymptotic decay rate and Bogoliubov instabilities
This leads to the following expression:
@@
$\lambda_{\mathrm{Bog}}=\left\{\begin{array}[]{l}\gamma\text{ if }\Delta<0\\ \gamma-\frac{2}{\sqrt{3}}\Delta\text{ if }0<\Delta<\frac{\sqrt{3}}{2}\gamma\\ 0\text{ else . }\end{array}\right.$ (S.4)
@@
This dependence is shown as the dotted line in Fig. S.4, alongside numerical results for the 2D square geometry. Both the analytical predictions and the numerical data display a step-like behavior with a linear slope between the upper and lower values. The two curves coincide for $\Delta\gtrsim 1.0\gamma$, where the system becomes critical and $\lambda_{\min}=0$. Deviations at other detuning values can be primarily attributed to finite-size effects.
@@
Figure S.4: Minimal asymptotic decay rate for the 2D square geometry (green squares), computed for a system size $L=81l$. Vertical error bars indicate the difference in $\lambda_{\min}$ between this size and a smaller system with $L=27l$. The grey dash-dotted line shows the linewidth $\lambda_{\mathrm{Bog}}$ predicted by Bogoliubov analysis for a uniform 2D system. The two curves exhibit similar behavior, with discrepancies attributed to finite-size effects.
VI Time evolution of the density profile

### v2

@@ Abstract
Our results were obtained by solving the multi-mode mean-field dynamics and analyzing the emergence of critical slowing down as the size of the driving patterns increases, and are supported by an analytical statistical mean-field theory that captures the dependence of the critical detuning on the Hausdorff dimension. An open and important question concerns the fate of the lower critical dimension in the presence of strong quantum boson-boson interactions, beyond the many-mode mean-field regime, particularly in the hard-core boson limit. Three scenarios could take place: (i) the lower critical dimension is the same in the strongly correlated regime; (ii) the lower critical dimension could evolve continuously with interaction strength, or (iii) a discontinuous shift could occur, signaling the emergence of a quantum phase transition with new phases in the strongly quantum correlated regime. Exploring this limit remains a key direction for future research. Another promising avenue is to investigate how fractional dimension and driven-dissipative conditions influence other classes of non-equilibrium phase transitions. This could yield deeper insight into the fundamental role of spatial geometry and dissipation in shaping critical phenomena in classical and quantum open systems.
@@ Acknowledgements.
“Tunable lower critical fractal dimension for a non-equilibrium phase transition”
@@
I Numerical method
We numerically simulate the driven–dissipative Gross–Pitaevskii equation [Eq. (1)] on a two-dimensional spatial grid. The system is initialized in the vacuum state, $\psi(\mathbf{r},t=0)=0$ for all $\mathbf{r}$. At time $t=0$, the coherent pump is switched on with a fixed driving amplitude $F$, corresponding to a driving intensity $I=F^{2}$. The equation is then time evolved up to a final time $t_{f}$, chosen sufficiently large for the dynamics to converge to the steady state. Time integration is performed using a second-order Runge–Kutta scheme with a fixed time step selected to ensure numerical stability.
To characterize the long-time dynamics, we compute the spatially averaged density $\bar{n}(t)=\int\mathrm{d}^{2}x\,n(\mathbf{r},t)$ and apply a Fourier transform to its time evolution. From the resulting spectrum, we extract the asymptotic decay rate $\lambda$ by measuring the full width at half maximum (FWHM) of the central peak, which provides the slowest relaxation rate of the system.
We have verified the robustness of our results by including quantum fluctuations within the Truncated Wigner Approximation, as shown in Fig. S1. The driven-dissipative Gross–Pitaevskii equation is found to capture the behavior remarkably well, as the spatial inhomogeneities induced by the driving pattern effectively act as spatial fluctuations [20].
Figure S1: First-order phase transition observed within the Truncated Wigner Approximation. Here, we consider the same 1D system as in the main text, with a longitudinal size $L=34.5\xi$ and $\Delta=3\gamma$. The observables are averaged over 100 trajectories.
II The effect of detuning for a Sierpinski carpet
@@
In the main text, we characterize critical behavior in terms of the minimal asymptotic decay rate. Here, we further examine how the asymptotic decay rate depends on the driving intensity for different system sizes. In Fig. S2(a)–(b), these quantities are shown for the Sierpinski carpet (second column in Fig. 1) with $\Delta=\gamma$, a regime where no phase transition occurs. The minimal asymptotic decay rate exhibits a clear minimum around a critical driving intensity $I\approx 2.2\gamma^{2}$, yet this minimum saturates to a finite value.
@@
Figure S2: Response of the system to a Sierpinski carpet drive with fractal dimension $d_{H}=1.84$. (a) Asymptotic decay rate as a function of input intensity for detuning $\Delta=\gamma$, showing no phase transition. (b) Corresponding bosonic density as a function of input intensity, with a smooth increase. (c) For a larger detuning $\Delta=2.0\gamma$, the minimal asymptotic decay rate $\lambda_{\min}$ vanishes at the critical point $I_{c}\approx 7.1\gamma^{2}$, indicating a first-order phase transition. (d) At the same critical point, the average bosonic density exhibits a discontinuous jump.
@@
In Fig. S2(c)–(d), we plot the same quantities for a larger detuning, $\Delta=2\gamma$. In this case, at a higher driving intensity $I_{c}\approx 7.1\gamma^{2}$, the minimal asymptotic decay rate $\lambda_{\min}$ collapses to numerical precision, and the density exhibits an abrupt switch from the lower to the upper branch of the bistability. These features indicate the occurrence of a first-order phase transition.
III Fractal dimension : nominal dimension vs box-counting dimension
In this section, we discuss different notions of fractal dimension and clarify the distinction between nominal and box-counting definitions. The nominal mathematical fractal dimension $d$ of a Sierpinski-carpet-like object can be determined from its self-similar scaling properties. If the pattern is constructed by dividing a square into $s\times s$ smaller squares and removing a given number of them (the “holes”), then at each iteration the remaining area scales as
$A=s^{2}-N_{\text{holes}}.$
This area corresponds to the white surface shown in Fig. S3. The nominal fractal dimension is then obtained from the ratio of logarithms,
$d^{th}_{H}=\frac{\log(A)}{\log(s)}=\frac{\log(s^{2}-N_{\text{holes}})}{\log(s)}.$
This expression characterizes how the number of occupied sites, or equivalently the total area, scales with the rescaling factor $s$ for a theoretical infinitely self-similar fractal. The procedure is illustrated in Fig. S3 for the specific case of the Sierpinski carpet.
simple square1st iteration2nd iteration
Figure S3: Constructing the Sierpinski Carpet. We divide an initial square in a grid of $3\times 3$ squares. We remove the central one. Then we can repeat this for the remaining eight squares. The area after one iteration is $A=s^{2}-N_{\text{holes}}=3^{2}-1=8$. So the mathematical theoretical dimension is $d^{th}_{H}=\log(8)/\log(3)=1.89$.
In a purely mathematical context, two different fractals can share exactly the same fractal dimension. This occurs, for example, when one chooses the same subdivision factor, $s$ and the same number of removed sites $N_{\text{holes}}$, but removes them in a different geometric arrangement at each iteration. One could follow the same construction procedure as in Fig. S3, but instead remove, say, the upper-left square rather than the central one; the resulting object would have a drastically different visual appearance (see Fig. S4 (left)), yet its nominal dimension $d^{th}_{H}$ would remain identical because the scaling rule encoded in $\log(A)/\log(s)$ is unchanged.
However, it is crucial to understand that in our physical system the final spatial profile of the bosonic density does not necessarily inherit this mathematical invariance: distinct driving geometries can lead to patterns that exhibit different effective fractal dimensions. This arises because the pattern is subject to physical distortions—such as nonlinear interactions, finite-resolution effects, and the intrinsic dynamics of the driven–dissipative medium—that modify how the nominal fractal mask is mapped onto the actual stationary density distribution. As a result, two drivings with mathematically identical fractal dimensions may generate bosonic patterns with noticeably different scaling behavior and thus different measured fractal dimensions.
@@ III Fractal dimension : nominal dimension vs box-counting dimension
To quantify the fractal dimension of the considered bosonic patterns, we employed the box-counting method [29]. The shape of each pattern is first extracted from a grayscale image (see, for example, Fig. 1) by applying a threshold above which points are considered part of the pattern [25]. This thresholding procedure delineates the fractal structure within the image, effectively separating the foreground (pattern) from the background and yielding a binary representation. In this binary image, pixels above the threshold are marked as belonging to the pattern, while those below are discarded. This representation serves as the basis for further spatial analysis.
@@
The box-counting method can be seen an inverse approach with respect to the fractal construction presented before. First, one divides the obtained binary image into a grid of $s^{2}$ square boxes, each with edge length $b=L/s$, where $L$ is the linear size of the system. We count the number of boxes occupied by the underlying density pattern, thereby estimating the area $A$ (in units of $b^{2}$) of the fractal, which corresponds to the white region shown in Fig. 1. In principle, this area scales as $A\propto s^{d_{H}}$, where $d_{H}$ is the Hausdorff (fractal) dimension. The box-counting dimension is thus given by the slope of $\log A=f(\log s)$ obtained over different scales $s$. By performing a linear fit to this curve, we extract an estimate of $d_{H}$, and the corresponding uncertainty is reported in the main text as the error bar on the box-counting dimension. The patterns in Fig. 1 are annotated with their respective box-counting dimensions $d_{H}$, which differ slightly from the theoretical nominal values $d^{th}_{H}$ for perfect, infinitely self-similar fractals. For example, for the Sierpinski carpet (second column), $d^{th}_{H}=1.89$, and for the Sierpinski cross (third column), $d^{th}_{H}=1.46$. Nevertheless, even for these distorted and finite structures, we can determine a fractal dimension lying between 1 and 2.
@@
For patterns with fractal dimension equal to or lower than 1 (Fig. 1, last column), we used a slightly modified version of the box-counting method. We restricted the analysis to the horizontal dimension by fixing the vertical box size to the system size $L$, while the horizontal box length still scales as $b=L/s$. Instead of square boxes of size $b^{2}$, we thus used rectangular boxes of size $bL$. This allows for a consistent evaluation of fractal dimensions ranging from 0 to 2. For integer-dimensional patterns (1D and 2D), we did not apply the box-counting method, as their Hausdorff dimensions are trivially $d_{H}=1$ and $d_{H}=2$, respectively. Nonetheless, one could apply box-counting to the patterns shown in the first and fourth columns of Fig. 1; such an analysis would yield values close to 1 or 2, but never exactly. Indeed, precisely capturing integer dimensions via box-counting is challenging due to inherent estimation errors.
At this point, we want to remind the reader explicitly that the value of the Hausdorff dimension calculated according to the equation $d_{H}^{\mathrm{th}}=\log(s^{2}-N_{holes})/\log(s)$ and the value obtained by box-counting may differ from each other. Nevertheless, we emphasize that the fractal box-counting dimension $d_{H}$ is one of the key ingredients in characterizing the critical behavior of our system. For our conclusions, it is not essential whether two driving schemes share the same theoretical dimension $d_{H}^{\mathrm{th}}$; what matters is the Hausdorff dimension of the resulting bosonic density pattern, which is determined via box counting. In the main text, we focus on symmetric driving geometries for clarity. We expect that fractals with asymmetric driven regions will also exhibit behavior governed by their fractal dimension. To illustrate this point, we consider an asymmetric pattern constructed following the same iterative procedure as for the Sierpinski carpet, but with the upper-left square removed instead of the central one. The nominal fractal dimension is therefore identical to that of the Sierpinski carpet, since in both cases $s=3$ and $N_{\mathrm{holes}}=1$, yielding $d_{H}^{\mathrm{th}}=\log(9-1)/\log(3)=1.89$. However, the resulting fractal (see Fig. S4 (left)) is no longer symmetric with respect to the central vertical axis.
Figure S4: Left: The spatial appearance of a pattern with asymmetric driving. Its construction is the same as in Fig. S3, except now the upper left square is removed. Right: An edited version of Fig. 3 of the main text where the assymetric patterns (the one shown on the left) is brought to the fore with a distinct colour (purple).
We can test that pattern in the same way we did in the main text. We obtain the result presented on the right-hand side of Fig. S4. The resulting box-counting Hausdorff dimensions $d_{H}$ predicts the critical dynamical behavior. For the asymmetric pattern (purple dot), we have very similar results to the Sierpinski carpet (blue empty square), which was to be expected since their nominal dimensions $d^{th}_{H}$ are the same.
IV The spatial profiles of the studied fractals
@@
Figure S5: Left panels: 2D square geometry ($d_{H}=2$) and several variations of the Sierpinski carpet with fractal dimensions in the range $1<d_{H}<2$. Right panels: Variations of the Sierpinski cross ($1<d_{H}<2$) in the first two rows, followed by the 1D strip ($d_{H}=1$) in the third row. The final row shows variations of the Cantor set with $d_{H}<1$. All patterns are computed with $\Delta=\gamma$, and the driving amplitudes are tuned to their respective critical points. The geometries and system sizes correspond to those used in Fig. 3.
@@
Figure S5 shows the spatial profiles of the patterns analyzed in the main text (Fig. 3). In addition to the standard 2D and 1D reference patterns, we include several variations of the Sierpinski cross and carpet, as well as modified versions of the Cantor set with fractal dimension lower than one.
V Dependence of the critical driving on the detuning
@@
In Fig. S6, we extract the value of the driving intensity $I_{\min}$ such that $\lambda(I_{\min})=\lambda_{\min}$, for a range of detuning values. As a reminder, $\lambda(I)$ denotes the asymptotic decay rate of the observed pattern as a function of the driving intensity $I$, and $\lambda_{\min}$ is its minimum value. As discussed in the main text, this minimum signals the presence of critical behavior in the system. Plotting $I_{\min}$ as a function of detuning thus allows us to trace how the onset of criticality evolves under varying external conditions.
@@
Interestingly, the value of $I_{\min}$ closely matches the inflection point of the bistability curve derived from single-mode ($k=0$) mean-field solution for the homogeneous driving case. The inflection point corresponds to the value of the driving intensity at which the curvature of the response function (4) changes sign—specifically, where the second derivative of the driving intensity with respect to the density $n=|\psi|^{2}$ vanishes:
$\frac{d^{2}I}{dn^{2}}=0~.$ (S1)
@@
$I_{\min}=\frac{2\Delta}{3g}\left(\frac{\Delta^{2}}{9}+\frac{\gamma^{2}}{4}\right)~,$
which is shown as the grey dashed line in Fig. S6. The close agreement between the numerical values of $I_{\min}$ and the analytical inflection point provides a consistency check and further supports the interpretation of $\lambda_{\min}$ as a reliable marker of criticality.
Figure S6: Evolution of the critical point $I_{\min}$ as a function of $\Delta$. A quadratic dependence is observed, consistent with estimates based on the single-mode mean-field description of the bistability curve. The inflection point of the mean-field curve (grey line) closely approximates the critical point across all geometries considered.
VI Asymptotic decay rate and Bogoliubov instabilities
For a homogeneous driving, we can diagonalize the problem for fluctuations on top of the steady state and obtain a spectrum of the so-called dissipative Bogoliubov modes [8]. By analyzing the dispersion relation given in equation (S2), we can interpret the phase transition as the emergence of an unstable mode.
$\omega_{\mathrm{Bog}}(\mathbf{k})=\pm\left[\left(-\Delta+\frac{\hbar\mathbf{k}^{2}}{2m}+2gn\right)^{2}-\left(gn\right)^{2}\right]^{1/2}-i\frac{\gamma}{2}~.$ (S2)
@@ VI Asymptotic decay rate and Bogoliubov instabilities
After calculations, this leads to the following expression:
@@
$\lambda_{\mathrm{Bog}}=\left\{\begin{array}[]{l}\gamma\text{ if }\Delta<0\\ \gamma-\frac{2}{\sqrt{3}}\Delta\text{ if }0<\Delta<\frac{\sqrt{3}}{2}\gamma\\ 0\text{ else . }\end{array}\right$. (S4)
@@
This dependence is shown as the dotted line in Fig. S7, alongside numerical results for the 2D square geometry. Both the analytical predictions and the numerical data display a step-like behavior with a linear slope between the upper and lower values. The two curves coincide for $\Delta\gtrsim 1.0\gamma$, where the system becomes critical and $\lambda_{\min}=0$. Deviations at other detuning values can be primarily attributed to finite-size effects.
@@
Figure S7: Minimal asymptotic decay rate for the 2D square geometry (green squares), computed for a system size $L=81l$. Vertical error bars indicate the difference in $\lambda_{\min}$ between this size and a smaller system with $L=27l$. The grey dash-dotted line shows the linewidth $\lambda_{\mathrm{Bog}}$ predicted by Bogoliubov analysis for a uniform 2D system. The two curves exhibit similar behavior, with discrepancies attributed to finite-size effects.
VII Time evolution of the density profile